OpenStudy (anonymous):

Evaluate the expression (-x^3y^2) for x=10 and y=-1 A.-1,000,000 B.-600,000 C.1,000,000 D.600,000

OpenStudy (anonymous):

\((-10)^3=-1000,(-1)^2=1\) is a start

OpenStudy (anonymous):

in fact none of your answers are correct, is there a typo there?

OpenStudy (anonymous):

nope

OpenStudy (anonymous):

then the answer is \(-1,000\)

OpenStudy (lgbasallote):

hmmm

OpenStudy (lgbasallote):

careful there..

OpenStudy (lgbasallote):

lol just kidding. yes it's supposed to be -1,000 unless there's a syntax error

OpenStudy (anonymous):

Sorry its (-x^3y^2)^2

OpenStudy (lgbasallote):

maybe it's \[(-x^3 y^2)^2\]

OpenStudy (lgbasallote):

hah i knew it*

OpenStudy (lgbasallote):

\[(-10^3 (-1)^2 )^2 \implies (-1000 \times 1)^2 \implies \cdots\] what do you think?

OpenStudy (anonymous):

nice thanks for the help

OpenStudy (lgbasallote):

<tips hat>

OpenStudy (anonymous):

ah an invisible square! good eye and carnak award for @lgbasallote